What is the electric field vector of a load that is at the origin of a coordinate axis at a distance of 50cm to a certain point P and makes an angle of 145 with respect to the axis of positive x and if the load of this is equal to Q = 8.9?
1
Expert's answer
2019-06-30T18:52:26-0400

The expression for determining the electric field vector at a distance "r" from the charge is


"E=\frac{1}{4\pi\epsilon_0}\cdot\frac{Q}{r^2}=\frac{1}{4\cdot3.14\cdot8.85\cdot10^{-12}}\cdot\frac{8.9}{0.5^2}=3.20\cdot10^{11}\text{ V/m}."